OpenStudy (anonymous):

Find three consecutive even integers such that the first times the third is 2 greater the 5 times than the second?

OpenStudy (unklerhaukus):

\[a\times(a+4)+2=5(a+2)\]

OpenStudy (anonymous):

let the even number be x......(x),(x+2),(x+4) are consecutive even numbers first times third =(x)(x+4) 5 times second is =(x+2)*5 according to the condition (x)(x+4)+2=5(x+2) re=read the condition and u'll understand....solve for x

OpenStudy (anonymous):

The Prophet and UnkleRhaukus thank you very much. :D

OpenStudy (unklerhaukus):

the first of the consecutive numbers is \(x \text{ or } a\) the second is \(x +2\text{ or } a+2\) the third is...

OpenStudy (anonymous):

I got it. x(x+4)=2+5(x+2) =x^2+4x=2+5x+10 =x^2+4x=12+5x =x^2-x-12=0 =(x-4)(x+3)

OpenStudy (unklerhaukus):

ah, so there are two sets of solutions, only one of these makes a sensible answer